A wheel of mass100kg and radius of gyration is 2m has a constant acceleration of 3rad/ sec² during a certain 4sec interval it turns through an angle of 120°rad.Assuming that wheel started from rest.How long was it in motion at the start of this 4sec interval.Find torque on wheel,kinetic energy during 4sec interval.

Grade:11

1 Answers

Vikas TU
7 years ago
From
2pi/3 = 0 + 0.5*3*t^2
we get,
t = 2root(pi)/3 seconds.
 
Torque during 4 seconds interval we get,
T = I*a = > 2*100*3 =.> 600 Kgm^2rad/s^2.
 
K.E = 0.5*I*w^2 = > 0.5*(pi/6)^2*200 = > 27.38 KJ